OpenStudy (adunb8):

undetermined coefficients... y"-y'+2y=-6xe^-x

OpenStudy (adunb8):

i know how to do make y = D \[D^2-D-2 = 0\] then \[(D+1)(D-2) \] there for it becomes \[C1e^-1x + C2e^2x\]

OpenStudy (adunb8):

for the general solution

OpenStudy (sirm3d):

particular solution first. \[\huge Axe^{-x}\]

OpenStudy (adunb8):

wher did Axe^-x come from?

OpenStudy (sirm3d):

multiple roots {-1}

OpenStudy (adunb8):

what?

OpenStudy (sirm3d):

my mistake. the particular solution is \[\large Axe^{-x} + Bx^2e^{-x}\]now you're really confused.

OpenStudy (adunb8):

yes i am can you show me how that happened?

OpenStudy (sirm3d):

let's begin with this. do you know ho to write a homogeneous DE for a particular solution, say \[\large y=5e^{3x}\]

OpenStudy (adunb8):

um.. not really 100% sure

OpenStudy (sirm3d):

let's do it backwards then. you concluded that the compl. solution to \[\large (D+1)(D-2)y = 0 \] is \[\large c_1e^{-x} + c_2e^{2x}\] is based on

OpenStudy (sirm3d):

is based on what result?

OpenStudy (adunb8):

general solution?

OpenStudy (sirm3d):

how did you get \[\huge e^{-x}\] and \[\huge e^{2x}\]? specifically, why did you choose -1 and +2 as coefficients of x?

OpenStudy (adunb8):

i found out that it was \[D^2-D-2 \] then factor it out \[(D+1)(D-2) \] which becomes D = -1, 2

OpenStudy (sirm3d):

we build our solution from that. \[\large e^{-x}\] for (D+1) and \[\large e^{2x}\] for (D-2) so for \[\large c_1e^{5x} + c_2e^{4x} + c_3e^{-2x} \] the polynomial in D is \[(D-5)(D-4)(D+2)\] agreed?

OpenStudy (adunb8):

what i dont understand is how did we get e^5x and e^4x?

OpenStudy (sirm3d):

\[\large (D-5)\] creates the solution \[\large ce^{5x}\] the coefficient of x is the root of (D-5). see the connection?

OpenStudy (adunb8):

yes i know that d-5 is ce^5x but where did d-5 come from that is the question

OpenStudy (sirm3d):

its not where it came from. when you see a (D-5) in a problem, you create \[\large ce^{5x}\] as part of the compl. solution to that problem where there is some (D-5)

OpenStudy (adunb8):

oh i see! thank you so much now i understand! THANKS for hanging with me!

OpenStudy (sirm3d):

are you still game?

OpenStudy (adunb8):

yea i actually understood it danka!

OpenStudy (sirm3d):

if you see (D-5)(D-5) in some problem, what functions do you create as part of the compl. solution?

OpenStudy (adunb8):

(D-5)^2 are you asking about that?

OpenStudy (sirm3d):

yes.